OpenStudy (avanti):

Which of the following is NOT an identity? A.1+cos^2x/sin^2x=1 B.(cosx + sinx)^2 = 1 + 2cosx sinx C(tanx + cotx)^2 = csc^2 x + sec^2 x D.tan^2x = sec^2x - sin^2x - cos^2x

OpenStudy (psymon):

Well, let's start with the first one and see what we can do with it. \[\frac{ 1+\cos ^{2}x }{ \sin ^{2}x }\] \[\frac{ 1 }{ \sin ^{2}x } + \frac{ \cos ^{2}x }{ \sin ^{2}x }\] \[\csc ^{2}x + \cot ^{2}x = 1\] Is this true? :P

OpenStudy (avanti):

no csc^2x - cot^2x=1

OpenStudy (psymon):

Well there ya go. Lucky us, first one was our answer.

OpenStudy (avanti):

okay thank you! :)

OpenStudy (psymon):

No problem ^_^

OpenStudy (avanti):

@Psymon can you help me with this one too please? which of the following IS an identity

OpenStudy (psymon):

Yeah, ill take a look.

OpenStudy (psymon):

Personally, I would start with the problems that actually look like something we could use. C and D look the simplest to mess around with, so I'll start with D

OpenStudy (psymon):

\[\sin ^{2}xcot ^{2}x + \cos ^{2}xtan ^{2}x = 1\] \[\frac{ \sin ^{2}xcos ^{2}x }{ \sin ^{2}x } + \frac{ \cos ^{2}xsin ^{2}x }{ \cos ^{2}x } = 1\] \[\cos ^{2}x + \sin ^{2}x = 1\] Lucky again, first one I tried :P. Really, you just gotta mess with the ones that actually look like something. I look at the first one and feel like I could mess with that forever and get nowhere, so I'd only do it as a last resort. 2nd one looks complex, too, so I just figured Id go in reverse order, do 4, 3, 2, 1 and hey, first one I tried :P

OpenStudy (avanti):

okay that makes sense, thank you so much! :D

OpenStudy (psymon):

Np :3